# On Hardy-Littlewood-Sobolev estimates for degenerate Laplacians

### v1

@@ 3.1. Key lemmas
where $P_{N}$ is a polynomial of degree $N-1$ with $P_{N}(0)=1$. Therefore, $\varphi_{\ell}(-\Delta_{\omega})=-P_{N}(\ell\Delta_{\omega})e^{\ell\Delta_{\omega}}+P_{N}(\ell^{-1}\Delta_{\omega})e^{{\ell}^{-1}\Delta_{\omega}}$. By Proposition 2.5, we have $\varphi_{\ell}(-\Delta_{\omega})f\to f$ in $L^{p}_{\omega}(\mathbb{R}^{n})$ as $\ell\to\infty$. Thus, $D((-\Delta_{\omega})^{-\alpha})\cap L^{p}_{\omega}(\mathbb{R}^{n})$ is dense in $L^{2}_{\omega}(\mathbb{R}^{n})\cap L^{p}_{\omega}(\mathbb{R}^{n})$, and hence dense in $L^{p}_{\omega}(\mathbb{R}^{n})$.                        $\square$
@@ 3.4. Extension to the functional calculus
We propose here a consequence of Theorem 1.1.
@@
Let $p\in(1,\infty)$. We say that $-\Delta_{\omega}$ admits a bounded holomorphic functional calculus on $L^{p}_{\omega}(\mathbb{R}^{n})$ if, for every $\mu\in(0,\pi)$ and every $\varphi\in\mathcal{H}^{\infty}_{0}(S_{\mu})$, and for all $f\in L^{2}_{\omega}(\mathbb{R}^{n})\cap L^{p}_{\omega}(\mathbb{R}^{n})$, the following estimate holds:
(3.12) $\|\varphi(-\Delta_{\omega})f\|_{L^{p}_{\omega}(\mathbb{R}^{n})}\leq C\|\varphi\|_{\infty,S_{\mu}}\|f\|_{L^{p}_{\omega}(\mathbb{R}^{n})},$
where $C>0$ is a constant independent of both $\varphi$ and $f$. By a standard density argument, inequality (3.12) implies that $\varphi(-\Delta_{\omega})$ extends to a bounded operator on $L^{p}_{\omega}(\mathbb{R}^{n})$. Moreover, this estimate remains valid for any bounded holomorphic function $\varphi$ on $S_{\mu}$. This extension follows readily by combining Fatou’s lemma with McIntosh’s convergence lemma in $L^{2}_{\omega}(\mathbb{R}^{n})$ (see [McI86]).
By combining the estimates in (2.5) with the fact that $-\Delta_{\omega}$ clearly has a bounded $H^{\infty}$-calculus on $L^{2}_{\omega}(\mathbb{R}^{n})$, one can apply a result due to Duong and Robinson [DR96, Theorem 3.1] concerning the boundedness of the $H^{\infty}$-calculus of $-\Delta_{\omega}$ on $L^{p}_{\omega}(\mathbb{R}^{n})$.
@@
The operator $-\Delta_{\omega}$ admits a bounded $H^{\infty}$-calculus on $L^{p}_{\omega}(\mathbb{R}^{n})$ for each $p\in(1,\infty)$.
@@
For the first point, we employ the standard argument (see [EHMT24, Theorem 5.15]) stating that if $f\in D((-\Delta_{\omega})^{\alpha})$ and $\psi$ is bounded, then $g=\psi(-\Delta_{\omega})f\in D((-\Delta_{\omega})^{\alpha})$, with $(-\Delta_{\omega})^{\alpha}g=\psi(-\Delta_{\omega})(-\Delta_{\omega})^{\alpha}f=\varphi(-\Delta_{\omega})f$. Applying inequality (1.1) to $g$ and substituting completes the argument.
@@
$\|\omega^{1/p}\psi(-\Delta_{\omega})f\|_{L^{q}_{\mathrm{d}x}(\mathbb{R}^{n})}=\|\omega^{1/p}(-\Delta_{\omega})^{-\alpha}g\|_{L^{q}_{\mathrm{d}x}(\mathbb{R}^{n})}\leq C\|g\|_{L^{p}_{\omega}(\mathbb{R}^{n})}\leq C^{\prime}\|\varphi\|_{\infty,S_{\mu}}\|f\|_{L^{p}_{\omega}(\mathbb{R}^{n})}.$
@@
By the same interpolation argument as in the proof of Theorem 3.4, when $\varphi$ is bounded, the operator $\psi(-\Delta_{\omega})$ is bounded from $L^{p}_{\omega}(\mathbb{R}^{n})$ to $L^{q,p}_{\omega}\big{(}\omega^{\frac{1}{p}-\frac{1}{q}},\mathbb{R}^{n}\big{)}$, with a bound of the form $C\|\varphi\|_{\infty,\mu}$, where $C$ is a positive constant depending only on $[\omega]_{A_{2}}$, $[\omega]_{RH_{\frac{q}{p}}}$, $n$, $p$, $q$, and $\mu$.
@@ 3.5. Sharpness of the condition on the weight
where $\tilde{C}<\infty$ is a positive constant depending only on $C$, $\alpha$, $p$, $q$, $n$, $D$ and the choice of $\mu$. Therefore, $\omega\in RH_{\frac{q}{p}}(\mathbb{R}^{n})$. ∎
@@ 4. Comparison with related results
For $\alpha=1/2$, then $I_{1}$ can be used to, roughly speaking, invert of the gradient and by construction $\|\nabla f\|_{L^{2}_{\omega}(\mathbb{R}^{n})}=\|(-\Delta_{\omega})^{1/2}f\|_{L^{2}_{\omega}(\mathbb{R}^{n})}$ on $H^{1}_{\omega}(\mathbb{R}^{n})$. So, we may also make use of the following result concerning the boundedness of the Riesz transform, originally established by Cruz-Uribe, Martell, and Rios [CUMR18].
Theorem 4.2 ([CUMR18]).
@@
Using these results, we see that one can obtain similar inequalities, only for $20<\alpha\leq 1/$ 2, in a smaller range of $p$ and assuming the weight in a smaller Muckenhoupt class and in dimension $n\geq 3$.
@@
In particular, $(-\Delta_{\omega})^{-\alpha/2}$ extends to a bounded operator from $L^{p}_{\omega}(\mathbb{R}^{n})$ to $L^{q}_{\omega^{q/p}}(\mathbb{R}^{n})$. Furthermore, if $p\in[2,\min(n,q_{+}(-\Delta_{\omega}))$ then same result holds, provided that $\omega\in A_{2}(\mathbb{R}^{n})\cap RH_{\frac{n}{n-p}}(\mathbb{R}^{n})$ and the constant $C$ depends only on $n$, $[\omega]_{A_{2}}$, $[\omega]_{RH_{\frac{n}{n-p}}}$ and $n$.
@@
The second condition is used qualitatively as follows. We have the pointwise inequality $|f|\leq C(n)|I_{1}(\nabla f)|$ for all $f\in\mathcal{D}(\mathbb{R}^{n})$. Using the density of $\mathcal{D}(\mathbb{R}^{n})$ in $H^{1}_{\omega}(\mathbb{R}^{n})$ together with Theorem 4.1 with $\alpha=1$, p=2 and $\tilde{\omega}=\omega^{\frac{1}{2}}$, we deduce that for all $f\in H^{1}_{\omega}(\mathbb{R}^{n})$, $f$ and $I_{1}(\nabla f)$ are almost everywhere limits of sequences $f_{k}$ and $I_{1}(\nabla f_{k})$ with $f_{k}\in\mathcal{D}(\mathbb{R}^{n})$. Hence $|f|\leq C(n)|I_{1}(\nabla f)|$ almost everywhere on $\mathbb{R}^{n}$, for all $f\in D((-\Delta_{\omega})^{1/2})=H^{1}_{\omega}(\mathbb{R}^{n})$.
@@
$\left\|f\right\|_{L^{p^{\star}}_{\omega^{p^{\star}/p}}(\mathbb{R}^{n})}\leq C(n)\left\|I_{1}(\nabla f)\right\|_{L^{p^{\star}}_{\omega^{p^{\star}/p}}(\mathbb{R}^{n})}\leq{C(n,p,[\omega]_{A_{p}},[\omega]_{RH_{\frac{n}{n-p}}})}\|\nabla f\|_{L^{p}_{\omega}(\mathbb{R}^{n})},\ \ \forall f\in D((-\Delta_{\omega})^{1/2}).$
Using Theorem 4.2, we deduce that for all $f\in D((-\Delta_{\omega})^{1/2})$,
@@
At this point, we apply an abstract version of Stein’s interpolation theorem due to Voigt [Voi92] to the analytic family $\Psi(z):=(-\Delta_{\omega})^{-\frac{z}{2}}$ defined for $0\leq\mathrm{Re}(z)\leq 1$, proceeding as follows. By the boundedness of the $H^{\infty}$-calculus on $L^{p}_{\omega}(\mathbb{R}^{n})$ in Lemma 3.9, $-\Delta_{\omega}$ has bounded imaginary powers. This, together with inequality (4.3) give us that for all $f\in D((-\Delta_{\omega})^{-1/2})\cap L^{p}_{\omega}(\mathbb{R}^{n})$,
@@
where $c_{0}<\infty$ and $c_{1}<\infty$ are finite constants that are independent of $f$ and that depend only on $n$, $p$, $[\omega]_{A_{p}}$ and $[\omega]_{RH_{\frac{n}{n-p}}}$. Using this abstract version of Stein’s interpolation theorem [Voi92], for all $\theta\in(0,1)$ and $f\in D((-\Delta_{\omega})^{-1/2})\cap L^{p}_{\omega}(\mathbb{R}^{n})$, we have
@@
Choosing $\theta\in(0,1)$ such that $\frac{1}{q}=\frac{1-\theta}{p}+\frac{\theta}{p^{\star}}$ and using the Stein-Weiss complex interpolation theorem with a change of measure (see [BL76, Theorem 5.5.3]), we have $[L^{p}_{\omega}(\mathbb{R}^{n}),L^{p^{\star}}_{\omega^{p^{\star}/p}}(\mathbb{R}^{n})]_{\theta}=L^{q}_{{\omega}^{q/p}}(\mathbb{R}^{n}),$ and we obtain
@@
We could have used [LMPT10, Theorem 2.7] instead of Theorem 4.1. The former provides a weighted Sobolev-Gagliardo-Nirenberg inequality involving the gradient, and it also includes the endpoint case $p=1$. This inequality is also sharp in the control of the bound with respect to the weight. More precisely, [LMPT10, Theorem 2.7] establishes that for any $p\geq 1$ and any weight $\tilde{\omega}\in A_{p,q}(\mathbb{R}^{n})$, where $q<\infty$ is such that $\frac{1}{p}-\frac{1}{q}=\frac{1}{n}$, the following estimate holds for every Lipschitz function $f$ with compact support:
@@
Remark that one must assume $n\geq 3$ to use it when $p=2$. In another direction, one can use [CUMR18, Proposition 6.1] which provides a bound $\|(-\Delta_{\omega})^{1/2}f\|_{L^{p}_{\omega}(\mathbb{R}^{n})}\lesssim\|\nabla f\|_{L^{p}_{\omega}(\mathbb{R}^{n})}$ when $p\in(r_{\omega},\infty)$ where $r_{\omega}<2$ is the infimum of exponents $r\geq 1$ for which $w\in A_{r}(\mathbb{R}^{n})$. Our result, together with this inequality, implies the above inequality in this range of $p$ and no information on the control of the constant.
@@ 5. Case of degenerate elliptic operators with coefficients
When the coefficients of $A$ are real-valued, all the previously stated results concerning the degenerate Laplacian $-\Delta_{\omega}$, such as the Gaussian upper (and lower) bounds for the heat kernel and the boundedness of the $H^{\infty}$-calculus on $L^{p}_{\omega}(\mathbb{R}^{n})$ for every $p\in(1,\infty)$ with the same references (see [AN25, Baa25] and [DR96, Theorem 3.1], respectively) remain valid for the degenerate second-order operator $L_{\omega}$, with the only modification being the sectoriality angle, which changes from 00 to $\varphi$. We may now adapts our results by replacing $-\Delta_{\omega}$ with $L_{\omega}$. As an illustration, the analogs of our results in Theorem 3.4 and Theorem 3.6 that just need the Gaussian upper bounds hold.
@@
Remark 5.2.
When the coefficients are complex, the reference [CUMR18] contains a thorough description of the properties of $L_{\omega}$. The only thing that changes for us is the existence of an interval $(p_{-}(L_{\omega}),p_{+}(L_{\omega}))\subset(1,\infty)$ as the largest open interval containing $2$ on which the semigroup $e^{-tL_{\omega}}$ is uniformly bounded on $L^{p}_{\omega}(\mathbb{R}^{n})$ for $p$ in this interval. One has $p_{-}(-\Delta_{\omega})=1$, $p_{+}(-\Delta_{\omega})=\infty$. The Gaussian upper bounds are replaced by a notion called $L^{p}_{\omega}-L^{q}_{\omega}$ off-diagonal estimates on balls (and balls could be equivalently replaced by cubes) denoted by $\mathcal{O}(L^{p}_{\omega}-L^{q}_{\omega})$. See Definition 2.33 and Proposition 3.1 in this reference, showing that they hold for $p<q$ in this interval. We claim that the proof of our key Lemma 2.3 can be adapted to obtain the same result as the first part of Theorem 5.1 but limited to $p_{-}(L_{\omega})<p<q<p_{+}(L_{\omega})$ (no weak type estimates at the endpoint). We shall leave details to the reader.
Remark 5.3.
Again, the results stated in Remark 5.2 above are partially obtainable from available ones from [MW74, LMPT10] at the expense of complications and restrictions on the hypotheses. For example, one can reproduce the argument in Corollary 4.3 to obtain (4.2) with $-\Delta_{\omega}$ replaced by a general degenerate elliptic operator $L_{\omega}$. However, one must first rely on the solution to the Kato problem [CUR15] for $L_{\omega}$ and generalisation [CUMR18] before applying the interpolation argument that follows.

### v2

@@ 3.1. Key lemmas
where $P_{N}$ is a polynomial of degree N1N-1 with $P_{N}(0)=1$. Therefore, $\varphi_{\ell}(-\Delta_{\omega})=-P_{N}(\ell\Delta_{\omega})e^{\ell\Delta_{\omega}}+P_{N}(\ell^{-1}\Delta_{\omega})e^{{\ell}^{-1}\Delta_{\omega}}$. By Proposition 2.10, we have $\varphi_{\ell}(-\Delta_{\omega})f\to f$ in $L^{p}_{\omega}(\mathbb{R}^{n})$ as $\ell\to\infty$. Thus, $D((-\Delta_{\omega})^{-\alpha})\cap L^{p}_{\omega}(\mathbb{R}^{n})$ is dense in $L^{2}_{\omega}(\mathbb{R}^{n})\cap L^{p}_{\omega}(\mathbb{R}^{n})$, and hence dense in $L^{p}_{\omega}(\mathbb{R}^{n})$.          $\square$
@@ 3.4. Extension to the functional calculus
We present here a consequence of Theorem 1.1.
@@
Let $p\in(1,\infty)$. We say that $\Delta_{\omega}$ admits a bounded holomorphic functional calculus on $L^{p}_{\omega}(\mathbb{R}^{n})$ if, for every $\mu\in(0,\pi)$ and every $\varphi\in\mathcal{H}^{\infty}_{0}(S_{\mu})$, and for all $f\in L^{2}_{\omega}(\mathbb{R}^{n})\cap L^{p}_{\omega}(\mathbb{R}^{n})$, the following estimate holds:
(3.20) $\|\varphi(-\Delta_{\omega})f\|_{L^{p}_{\omega}(\mathbb{R}^{n})}\leq C\|\varphi\|_{\infty,\mu}\|f\|_{L^{p}_{\omega}(\mathbb{R}^{n})},$
where $C>0$ is a constant independent of both $\varphi$ and $f$, and where $\|\varphi\|_{\infty,\mu}$ denotes the $L^{\infty}$-norm of $\varphi$ on the sector $S_{\mu}$. By a standard density argument, inequality (3.20) implies that $\varphi(-\Delta_{\omega})$ extends to a bounded operator on $L^{p}_{\omega}(\mathbb{R}^{n})$. Moreover, this estimate remains valid for any bounded holomorphic function $\varphi$ on $S_{\mu}$. This extension follows readily by combining Fatou’s lemma with McIntosh’s convergence lemma in $L^{2}_{\omega}(\mathbb{R}^{n})$ (see [MCI86]).
By combining the estimates in (2.9) with the fact that $\Delta_{\omega}$ clearly has a bounded $H^{\infty}$-calculus on $L^{2}_{\omega}(\mathbb{R}^{n})$, one can apply a result due to Duong and Robinson [DR96, Theorem 3.1] concerning the boundedness of the $H^{\infty}$-calculus of $\Delta_{\omega}$ on $L^{p}_{\omega}(\mathbb{R}^{n})$.
@@
The operator $\Delta_{\omega}$ admits a bounded $H^{\infty}$-calculus on $L^{p}_{\omega}(\mathbb{R}^{n})$ for each $p\in(1,\infty)$.
@@
For the first point, we employ the standard argument (see [EHM+24, Theorem 5.15]) stating that if $f\in D((-\Delta_{\omega})^{\alpha})$ and $\psi$ is bounded, then $g=\psi(-\Delta_{\omega})f\in D((-\Delta_{\omega})^{\alpha})$, with $(-\Delta_{\omega})^{\alpha}g=\psi(-\Delta_{\omega})(-\Delta_{\omega})^{\alpha}f=\varphi(-\Delta_{\omega})f$. Applying inequality (1.2) to $g$ and substituting completes the argument.
@@
$\|\omega^{1/p}\psi(-\Delta_{\omega})f\|_{L^{q}_{\mathrm{d}x}(\mathbb{R}^{n})}=\|\omega^{1/p}(-\Delta_{\omega})^{-\alpha}g\|_{L^{q}_{\mathrm{d}x}(\mathbb{R}^{n})}\leq C\|g\|_{L^{p}_{\omega}(\mathbb{R}^{n})}\leq C^{\prime}\|\varphi\|_{\infty,\mu}\|f\|_{L^{p}_{\omega}(\mathbb{R}^{n})}.$
@@
By the same interpolation argument as in the proof of Theorem 3.15, when $\varphi$ is bounded, the operator $\psi(-\Delta_{\omega})$ is bounded from $L^{p}_{\omega}(\mathbb{R}^{n})$ to $L^{q,p}_{\omega}\big(\omega^{\frac{1}{p}-\frac{1}{q}},\mathbb{R}^{n}\big)$, with a bound of the form $C\|\varphi\|_{\infty,\mu}$, where $C$ is a positive constant depending only on $\omega]_{A_{2}}$, $\omega]_{RH_{\frac{q}{p}}}$, $n$, $p$, $q$, and $\mu$.
@@ 3.5. Sharpness of the condition on the weight
where $\tilde{C}$$\infty$ is a positive constant depending only on $C$, $\alpha$, $p$, $q$, $n$, $D$ and the choice of $\mu$. Therefore, $\omega\in RH_{\frac{q}{p}}(\mathbb{R}^{n})$. ∎
@@ 4. Comparison with related results
For $\alpha=1/2$, then $I_{1}$ can be used to, roughly speaking, invert of the gradient and by construction $\|\nabla f\|_{L^{2}_{\omega}(\mathbb{R}^{n})}=\|(-\Delta_{\omega})^{1/2}f\|_{L^{2}_{\omega}(\mathbb{R}^{n})}$ on $H^{1}_{\omega}(\mathbb{R}^{n})$. So, we may also make use of the following result concerning the boundedness of the Riesz transform, originally established by Cruz-Uribe, Martell, and Rios [CMR18].
Theorem 4.2 ([CMR18]).
@@
Using these results, we see that one can obtain similar inequalities, only for $0<\alpha\leq 1/2$, in a smaller range of $p$ and assuming the weight in a smaller Muckenhoupt class and in dimension $n\geq 3$.
@@
In particular, $(-\Delta_{\omega})^{-\alpha/2}$ extends to a bounded operator from $L^{p}_{\omega}(\mathbb{R}^{n})$ to $L^{q}_{\omega^{q/p}}(\mathbb{R}^{n})$. Furthermore, if $p\in[2,\min(n,q_{+}(-\Delta_{\omega}))$ then same result holds, provided that $\omega\in A_{2}(\mathbb{R}^{n})\cap RH_{\frac{n}{n-p}}(\mathbb{R}^{n})$ and the constant $C$ depends only on $n$, $\omega]_{A_{2}}$, $\omega]_{RH_{\frac{n}{n-p}}}$ and $n$.
@@
The second condition is used qualitatively as follows. We have the pointwise inequality $|f|\leq C(n)|I_{1}(\nabla f)|$ for all $f\in\mathcal{D}(\mathbb{R}^{n})$. Using the density of $\mathcal{D}(\mathbb{R}^{n})$ in $H^{1}_{\omega}(\mathbb{R}^{n})$ together with Theorem 4.1 with $\alpha=1$, p=2 and $\tilde{\omega}=\omega^{\frac{1}{2}}$, we deduce that for all $\in H^{1}_{\omega}(\mathbb{R}^{n}), ff$ and $I_{1}(\nabla f)$ are almost everywhere limits of sequences $f_{k}$ and $I_{1}(\nabla f_{k})$ with $f_{k}\in\mathcal{D}(\mathbb{R}^{n})$. Hence $|f|\leq C(n)|I_{1}(\nabla f)|$ almost everywhere on $\mathbb{R}^{n}$, for all $f\in D((-\Delta_{\omega})^{1/2})=H^{1}_{\omega}(\mathbb{R}^{n})$.
@@
$\left\|f\right\|_{L^{p^{\star}}_{\omega^{p^{\star}/p}}(\mathbb{R}^{n})}\leq C(n)\left\|I_{1}(\nabla f)\right\|_{L^{p^{\star}}_{\omega^{p^{\star}/p}}(\mathbb{R}^{n})}\leq{C(n,p,[\omega]_{A_{p}},[\omega]_{RH_{\frac{n}{n-p}}})}\|\nabla f\|_{L^{p}_{\omega}(\mathbb{R}^{n})},$
for all $f\in D((-\Delta_{\omega})^{1/2})$. Using Theorem 4.2, we deduce that for all $f\in D((-\Delta_{\omega})^{1/2})$,
@@
At this point, we apply an abstract version of Stein’s interpolation theorem due to Voigt [VOI92] to the analytic family $\Psi(z):=(-\Delta_{\omega})^{-\frac{z}{2}}$ defined for $0\leq\mathrm{Re}(z)\leq 1$, proceeding as follows. By the boundedness of the $H^{\infty}$-calculus on $L^{p}_{\omega}(\mathbb{R}^{n})$ in Lemma 3.21, $\Delta_{\omega}$ has bounded imaginary powers. This, together with inequality (4.6) give us that for all $f\in D((-\Delta_{\omega})^{-1/2})\cap L^{p}_{\omega}(\mathbb{R}^{n})$,
@@
where $c_{0}<\infty$ and $c_{1}<\infty$ are finite constants that are independent of $f$ and that depend only on $n$, $p$, $\omega]_{A_{p}}$ and $\omega]_{RH_{\frac{n}{n-p}}}$. Using this abstract version of Stein’s interpolation theorem [VOI92], for all $\theta\in(0,1)$ and $f\in D((-\Delta_{\omega})^{-1/2})\cap L^{p}_{\omega}(\mathbb{R}^{n})$, we have
@@
Choosing $\theta\in(0,1)$ such that $\frac{1}{q}=\frac{1-\theta}{p}+\frac{\theta}{p^{\star}}$ and using the Stein-Weiss complex interpolation theorem with a change of measure (see [BL76, Theorem 5.5.3]), we have $L^{p}_{\omega}(\mathbb{R}^{n}),L^{p^{\star}}_{\omega^{p^{\star}/p}}(\mathbb{R}^{n})]_{\theta}=L^{q}_{{\omega}^{q/p}}(\mathbb{R}^{n}),$ and we obtain
@@
We could have used [LMP+10, Theorem 2.7] instead of Theorem 4.1. The former provides a weighted Sobolev-Gagliardo-Nirenberg inequality involving the gradient, and it also includes the endpoint case $p=1$. This inequality is also sharp in the control of the bound with respect to the weight. More precisely, [LMP+10, Theorem 2.7] establishes that for any $p\geq 1$ and any weight $\tilde{\omega}\in A_{p,q}(\mathbb{R}^{n})$, where $q<\infty$ is such that $\frac{1}{p}-\frac{1}{q}=\frac{1}{n}$, the following estimate holds for every Lipschitz function $f$ with compact support:
@@
Remark that one must assume $n\geq 3$ to use it when $p=2$. In another direction, one can use [CMR18, Proposition 6.1] which provides a bound $\|(-\Delta_{\omega})^{1/2}f\|_{L^{p}_{\omega}(\mathbb{R}^{n})}\lesssim\|\nabla f\|_{L^{p}_{\omega}(\mathbb{R}^{n})}$ when $p\in(r_{\omega},\infty)$ where $r_{\omega}<2$ is the infimum of exponents $r\geq 1$ for which $w\in A_{r}(\mathbb{R}^{n})$. Our result, together with this inequality, implies the above inequality in this range of $p$ and no information on the control of the constant.
@@ 5. Case of degenerate elliptic operators with coefficients
When the coefficients of $A$ are real-valued, all the previously stated results concerning the degenerate Laplacian $\Delta_{\omega}$, such as the Gaussian upper (and lower) bounds for the heat kernel and the boundedness of the $H^{\infty}$-calculus on $L^{p}_{\omega}(\mathbb{R}^{n})$ for every $p\in(1,\infty)$ with the same references (see [AN25, BAA26] and [DR96, Theorem 3.1], respectively) remain valid for the degenerate second-order operator $L_{\omega}$, with the only modification being the sectoriality angle, which changes from 0 to $\varphi$. We may therefore adapt results that just need the Gaussian upper bounds to hold by replacing $\Delta_{\omega}$ with $L_{\omega}$. As an illustration, the analog of Theorem 3.15 and Theorem 3.17 is as follows.
@@
When the coefficients are complex, the reference [CMR18] provides us with a thorough description of the properties of $L_{\omega}$. In particular, it shows the existence of an interval $(p_{-}(L_{\omega}),p_{+}(L_{\omega}))\subset(1,\infty)$, which is the largest open interval containing $2$ of exponents $p$ for which the semigroup $e^{-tL_{\omega}}$ is uniformly bounded on $L^{p}_{\omega}(\mathbb{R}^{n})$. One has $p_{-}(-\Delta_{\omega})=1$ and $p_{+}(-\Delta_{\omega})=\infty$. The Gaussian upper bounds are replaced by a notion called $L^{p}_{\omega}-L^{q}_{\omega}$ off-diagonal estimates on balls (which can equivalently be replaced by cubes), denoted by $\mathcal{O}(L^{p}_{\omega}-L^{q}_{\omega})$. See Definition 2.33 and Proposition 3.1 in that reference. It is not clear, however, how to proceed in our setting with this notion. Moreover, there is a difference of point of view. We interpret weighted spaces via multiplication by the weight, which allows us to extend techniques from the unweighted theory. In particular, the target space is endowed with the measure $\omega^{q/p}\mathrm{d}x$. In contrast, in [CMR18], the results are formulated in the space of homogeneous type $(\mathbb{R}^{n},\mathrm{d}w,|\cdot|)$.
We mention that, even in the absence of Gaussian upper bounds, one may still obtain some results from [MW74, LMP+10], at the expense of strong hypotheses. More precisely, one can reproduce the argument in Corollary 4.3 to obtain (4.5) with $\Delta_{\omega}$ replaced by $L_{\omega}$, under the assumptions that $n\geq 3$, $p_{-}(L_{\omega})<p<q_{+}(L_{\omega})$ and $\omega\in A_{p}(\mathbb{R}^{n})\cap RH_{\frac{n}{n-2}}(\mathbb{R}^{n})$. In addition, one needs the solution to the Kato problem for $L_{\omega}$ in [CR15] and its $L^{p}_{\omega}$ generalization in [CMR18] before applying the subsequent interpolation argument.
Another interesting remark is that when $p=2$ and $\alpha\in(0,1/2)$, one can make use of the results just established for $(-\Delta_{\omega})^{\alpha}$. Indeed, Kato’s theorem [KAT61] yields $D(L_{\omega}^{\alpha})=D((-\Delta_{\omega})^{\alpha})$, and the norms $\|L_{\omega}^{\alpha}\cdot\|_{2,\omega}$ and $\|(-\Delta_{\omega})^{\alpha}\cdot\|_{2,\omega}$ are equivalent, independently of the coefficients.